# BertaQA: How Much Do Language Models Know About Local Culture?

### v1

@@ 2 BertaQA
The dataset was originally compiled in Basque by crawling public sources that are no longer available. Google does not return any result when searching for questions from the dataset verbatim. While this cannot categorically discard contamination, we believe that this, along with the nature of the raw data we crawled and the results from our experiments, makes it very unlikely that existing models were exposed to the same data during training.
Starting from the original version in Basque, we also created an English version of BertaQA using a professional translation service. Translators were instructed to use a consistent format for all the questions and answers, and we refined our guidelines through multiple rounds. For named entities, Wikipedia was used as a reference when available. During the translation process, a few of the original questions in Basque were corrected, either because the original answer was incorrect or it became outdated. In addition, we discarded a few questions that required knowledge of Basque or English, and would lose their essence if translated.
@@ 4.1 Main results
We first evaluate existing models in English, focusing on the difference in performance between local and global topics. As shown in Table 2, most models obtain good results on global questions. However, the performance consistently drops when evaluated on local topics, with a gap of 26.66 points on average. More concretely, state-of-the-art models like GPT-4 Turbo and Claude 3 Opus shine on the global subset, scoring above 90% accuracy. Nevertheless, these same models obtain about 72% accuracy on the local subset. The difference is even larger for open-weight models, with the best one (Llama 3 70B) obtaining less than 60% accuracy on local questions. This confirms that, despite the impressive performance of LLMs in knowledge-intensive tasks, subjects pertinent to minority cultures remain challenging. As reported in Appendices E and F, results further drop to around 53% for the worst category and 66% for the hardest difficulty, leaving ample room for improvement for future work.
Interestingly, we find that the performance on local and global questions is strongly correlated for the models we tested (i.e., models obtaining a similar score on the local subset also obtain a similar score on the global subset). This suggests that the training recipes of existing models are roughly equivalent in how they balance global and local knowledge. However, we do find notable differences on how scaling impacts local vs. global questions for different models. For weaker model families like Llama 2, scaling yields bigger gains on the global subset, as the delta between local and global questions increases from 22.80 for the smallest variant to 28.53 for the largest variant. The opposite is true for stronger model families like GPT, with the delta between local and global questions going from 27.32 for GPT-3.5 Turbo to 19.51 for GPT-4 Turbo. This suggests that it is generally easier to improve on global questions, but this subset starts saturating for the strongest models, resulting in bigger improvements on the local subset.
@@ 4.2 Local knowledge transfer from Basque to English
To test this hypothesis, we experiment with Latxa [Etxaniz et al., 2024], a family of Basque language models that were built by continuing training Llama 2 on Basque corpora. As shown in Table 3, this continued training in Basque brings large improvements on the local subset, outperforming the original Llama 2 model by 13.46 points in the case of the 70B variant. It is remarkable that we observe these gains when performing evaluation in English, even if the continued training is done in Basque, which implies that there is knowledge transfer from Basque into English. This challenges the conventional wisdom that adding more languages hurts English performance, a phenomenon known as the curse of multilinguality [Conneau et al., 2020, Pfeiffer et al., 2022, Chang et al., 2024]. To the best of our knowledge, this is the first solid evidence of knowledge being transferred from a low-resource to a high-resource language.
Nevertheless, we observe the opposite effect on the global subset, where the continued training in Basque hurts performance. The degradation is relatively small for the 13B and 70B models, but more notable for the 7B model. This suggests that training on Basque data improves English performance on subjects related to Basque culture, while hurting performance on more general topics. Given that prior work mostly evaluated on global subjects, this led to the incomplete conclusion that training on other languages harms English.
@@ 5 Related Work
Research in NLP evaluation has predominantly focused in English, with most multilingual benchmarks being translated from this language, such as XNLI [Conneau et al., 2018], XQUAD [Artetxe et al., 2019], MLQA [Lewis et al., 2019] and Belebele [Bandarkar et al., 2023]. This parallel nature facilitates monolingual, multilingual, and cross-lingual experiments, enabling valuable comparisons across languages. However, this approach introduces biases related to translations and cultural representation, affecting experimental conclusions by reflecting only the culture of the original English dataset.
Recently, there has been a focus on creating native evaluation benchmarks to assess local cultural knowledge, rather than relying on translations from English. These native datasets, which resemble popular English benchmarks, include unique cultural elements that are generally more challenging for current models. They usually are of higher quality than machine or human-translated datasets. For example, native MMLU [Hendrycks et al., 2020] datasets have been created for Chinese [Li et al., 2023], Korean [Son et al., 2024], Indonesian [Koto et al., 2023] and Arabic [Koto et al., 2024]. Other examples of language-specific evaluation benchmarks include C-Eval for Chinese [Huang et al., 2024], HAE-RAE Bench for Korean [Son et al., 2023] and COPAL-ID for Indonesian [Wibowo et al., 2023]. Finally, Etxaniz et al. [2024] introduces 4 native Basque multiple-choice evaluation datasets that include local questions.
@@ 6 Conclusion and Future Work
While our work is a first step in this direction, it also comes with some limitations, which we would like to address in future work. More concretely, the local subset of our benchmark is limited to questions about the Basque culture. While we expect that the general trends we observe would also apply to other minority cultures, we believe that it would be valuable to build similar datasets covering other cultures. This is generally more challenging than developing benchmarks about global subjects, as it usually requires being part of or engaging with the relevant communities. Besides, the English version of our dataset was created through professional translation, which could lead to translationese and other translation artifacts [Artetxe et al., 2020]. This typically occurs in the opposite direction, as most datasets are translated from English into other languages. In the future, we would like to analyze if this has any impact in our results, and explore authoring questions about minority cultures in English rather than translating from their respective languages.
@@ Appendix E Results by Category
We show that previous local and global results are consistent across categories in Tables 9 and 10. The large difference between local and global is maintained across all models and categories. When comparing Llama 2 and Latxa, we see that previous results are consistent across all categories. That is, Latxa is better at local questions and Llama 2 is better at global questions. However, we see that the differences vary significantly depending on the models and categories.
@@ Appendix F Results by Difficulty
Results are also consistent across difficulty levels, as shown in Tables 9 and 12. All models obtain worse results in local questions at all difficulty levels. There is a clear difference between difficulties, with the biggest drop in performance happening from the easy to medium difficulty. Commercial models have a bigger drop in scores on local questions than on global questions for medium and hard questions. For open models, results are more varied, but the relative drop is also generally bigger on local questions. On the most difficult local questions, some models get close to random chance.

### v2

@@ 2 BertaQA
The dataset was originally compiled in Basque by crawling public sources that are no longer available. The questions were already classified into local and global topics, and labeled according to their difficulty level and knowledge category. We inspected the dataset and confirmed that the division was well-founded. The motivation for using this global subset, as opposed to existing QA datasets, is that the questions come from the same source, so the results should be more comparable. As such, no human annotator was involved in the creation of the Basque portion of the dataset. To check whether the content is present in other websites, we wrote some of the questions verbatim in Google search using quotation marks, but received no results. Finding the answer to some of these questions on the web is still possible, but at least the same questions are not present on the web. We do this experiment in Section 4.5, where we try to find the correct answer of some questions by searching the web. While this cannot categorically discard contamination, we believe that this, along with the nature of the raw data we crawled and the results from our experiments, makes it very unlikely that existing models were exposed to the same data during training.
Starting from the original version in Basque, we also created an English version of BertaQA using a professional translation service (Elhuyar itzulpenak [^4] https://itzulpenak.elhuyar.eus/en). We first wrote some translation guidelines, covering things like formatting or using Wikipedia as a reference when available to translate Basque named entities. In addition, our guidelines asked translators to discard questions whose answers require knowing the Basque language (such as onomatopeias). We initially sent 100 question/answers for translation. We reviewed these translations carefully, and worked closely with the professional translators to clarify and extend the guidelines accordingly. The remaining dataset was translated in batches of 1000 samples. The translators tagged problematic samples (difficult translation, outdated information, more than one correct answer…), which we manually reviewed. During the translation process, a few of the original questions in Basque were corrected, either because the original answer was incorrect or it became outdated. In addition, we discarded a few questions that required knowledge of Basque or English, and would lose their essence if translated.
@@ 4.1 Main results
We first evaluate existing models in English, focusing on the difference in performance between local and global topics. As shown in Table 2, most models obtain good results on global questions. However, the performance consistently drops when evaluated on local topics, with a gap of 26.66 points on average. More concretely, state-of-the-art models like GPT-4 Turbo and Claude 3 Opus shine on the global subset, scoring above 90% accuracy. Nevertheless, these same models obtain about 72% accuracy on the local subset. The difference is even larger for open-weight models, with the best one (Llama 3 70B) obtaining less than 60% accuracy on local questions. This confirms that, despite the impressive performance of LLMs in knowledge-intensive tasks, subjects pertinent to minority cultures remain challenging. As reported in Appendices E and F, results further drop to around 53% for the worst category and 66% for the hardest difficulty, leaving ample room for improvement for future work. We manually inspected a subset of 50 local questions that the best models (GPT 4 Turbo and Claude 3 Opus) answered incorrectly in Section 4.5.
Interestingly, we find that the performance on local and global questions is strongly correlated for the models we tested (the Pearson correlation between the two scores is 0.844). Models obtaining a similar score on the local subset also obtain a similar score on the global subset. We presume that, if the training corpus of a given model was significantly more skewed towards local topics than that of another model, the former would tend to perform better in local topics, at least in relative terms. Given that we do not observe this, we hypothesize that the training recipes of existing models are roughly equivalent in how they balance global and local knowledge. However, we do find notable differences on how scaling impacts local vs. global questions for different models. For model families with the lowest scores in BertaQA, like Llama 2, scaling yields bigger gains on the global subset, as the delta between local and global questions increases from 22.80 for the smallest variant to 28.53 for the largest variant. The opposite is true for more performant model families like GPT, with the delta between local and global questions going from 27.32 for GPT-3.5 Turbo to 19.51 for GPT-4 Turbo. This suggests that it is generally easier to improve on global questions, but this subset starts saturating for the strongest models, resulting in bigger improvements on the local subset.
@@ 4.2 Local knowledge transfer from Basque to English
To test this hypothesis, we experiment with Latxa [Etxaniz et al., 2024], a family of Basque base language models that were built by continuing training Llama 2 on Basque corpora. Latxa was trained on all publicly available corpora in Basque meeting some minimum quality standards. This mostly corresponds to crawling corpora, including both processed versions of CommonCrawl as well as ad-hoc crawling of websites with high-quality content. The largest portion of it consists of news, which is the most common use of Basque on the web. While the resulting pretraining corpus is diverse in nature, we would not say that the domains and topics in BertaQA are particularly prominent on it, although it is obviously more likely that topics related to the Basque culture are discussed in the Basque language compared to English. As shown in Table 3, this continued training in Basque brings large improvements on the local subset, outperforming the original Llama 2 model by 13.46 points in the case of the 70B variant. It is remarkable that we observe these gains when performing evaluation in English, even if the continued training is done in Basque, which implies that there is knowledge transfer from Basque into English. This challenges the conventional wisdom that adding more languages hurts English performance, a phenomenon known as the curse of multilinguality [Conneau et al., 2020, Pfeiffer et al., 2022, Chang et al., 2024]. To the best of our knowledge, this is the first solid evidence of knowledge being transferred from a low-resource to a high-resource language.
Nevertheless, we observe the opposite effect on the global subset, where the continued training in Basque hurts performance. The degradation is relatively small for the 13B and 70B models, but more notable for the 7B model. This suggests that training on Basque data improves English performance on subjects related to Basque culture, while hurting performance on more general topics. Given that prior work mostly evaluated on global subjects, this led to the generally accepted conclusion that training on other languages harms English. We show that this conclusion does not necessarily show the full picture, since models have barely been evaluated on local topics, and their behavior there can be fundamentally different.
@@ 4.4 Translate-test and self-translate
4.5 Error Analysis
As we stated previously, some local questions are still challenging even for the best models. To get more insights of the nature of these questions, we do a qualitative analysis of 50 local questions in English that the best models (GPT 4 Turbo and Claude 3 Opus) answered incorrectly. As expected, most questions in this sample are of medium or hard difficulty, and fall in the most difficult categories. None of the questions can be answered by common sense, and distractors are generally challenging.
Next, we try to find the correct answer of these questions by searching the web, to measure how difficult they can be. We include half of the examples in Table 6, the rest can be found in Appendix G. First, we found the correct answer relatively easily in 24 of the questions. Next, in 20 questions, finding the correct answer was more challenging, requiring multiple web searches, or searching the web in Basque. Six of these answers were trickier to find in English, often leading to no answer or even incorrect answers. Finally, in 16 of the questions we were unable to find the correct answer. In half of the questions, searching the web led to the wrong answer. Some of these questions involve temporal variations, the correct answer has changed in time. There were also some questions where we found no answer, such as the ones including negation.
Table 6: First 25 error analysis examples annotated by difficulty of web search.
Category Diff. Question Candidate 0 Candidate 1 Candidate 2 Ans. Web Basque and Literature 1 What does the "Karmel" magazine specialize in? Bertsolarism Basque culture in the past and the present The life of the Carmelites 1 diff. Geography and History 3 Which of these towns does not have a common boundary with Elorrio? Zaldibar Bergara Matiena 2 easy Geography and History 3 Which of these districts is not in the Pettarra area of Soule? Etxarri Garruze Sarricotapea 1 no Sports and Leisure 2 What’s Aritz Aranburu’s birthplace? Zarautz Orio Getaria 2 easy Basque and Literature 2 Which Basque dialect is spoken in Eibar? The Lower
Deba dialect
The Central
dialect
The Western
dialect
2 easy
Basque and Literature 2 Which of these three was the Head of a Department in the Chartered Provincial Council of Gipuzkoa? Tere
Irastortza
Joan Mari
Irigoien
Xabier Lete 2 diff.
Cinema and Shows 2 The stories of how many young people is the "Hasiberriak" TV show about? Ten Seven Five 0 diff. Geography and History 1 Which is the longest river that flows into the sea in the Basque Country? The Nervion The Adour The Ebro 1 diff. Geography and History 2 Which of these is the least populated area? The canton of
Maule/Mauleón
The canton of
Donapaleu/
Saint-Palais
The canton of
Atharratze/
Tardets
1 diff.
Science and Technology 2 How many echidna species are there? 8 5 3 1 no Basque and Literature 1 What is Ricardo Arregi Diaz de Heredia mostly involved in? Journalism Adventure stories Poetry 2 easy Geography and History 3 What year was "La Vizcaya" factory set up? In 1882 In 1884 In 1886 0 no Geography and History 1 What century does the oldest Basque text we know written on paper belong to? The 15th century The 12th century The 10th century 2 no Music and Dance 1 Where was the great dance master Iñaki Irigoien from? From Bilbao From Donostia / San Sebastian From Vitoria-Gasteiz 0 easy Geography and History 3 Which king of Navarre died in 882? Fortun Gartzia Eneko Aritza Gartzia Eneko 2 difficult Cinema and Shows 2 Which band wrote the opening song of the "Pilotari" TV series? Gari Sugan Ken Zazpi 1 difficult Basque and Literature 2 Which of these subjects was not addressed at the Basque Floral Games? Dance Rural sports Basque poetry 1 no Sports and Leisure 3 When was the Leurtza reservoir built? In 1920 In 1925 In 1921 0 easy Science and Technology 2 Where’s the Basque Museum of Medical History? In Bilbao In Leioa In Barakaldo 1 easy Music and Dance 2 Where does the "Axuri Beltza" dance come from? From the Biscayan town of Aulestia From the Gipuzkoan town of Errezil From the Navarrese town of Jaurrieta 2 difficult Cinema and Shows 3 When is the "Teknopolis" programme broadcast on ETB1 (Basque Public TV)? From Monday to Thursday at 18:00 On Saturdays at 15:00 On Fridays and Saturdays at 11:00 1 no Science and Technology 2 What is the other name of the first blast furnace of Altos Hornos de Vizcaya? The Miren Agote furnace The Maria Angeles furnace The Santa Ana furnace 1 no Music and Dance 3 How many dance championships are held in the Northern Basque Country a year? None Three Five 0 no Music and Dance 2 Where does the "Trapatan" dance take place today? In Etxarri Aranatz In Ituren In Doneztebe/Santesteban 2 difficult Cinema and Shows 2 How many Basque voices took part in the show "Sortuko dira besteak"? Six Eight Four 0 difficult
@@ 5 Related Work
Research in NLP evaluation has predominantly focused in English, with most multilingual benchmarks being translated from this language, such as XNLI [Conneau et al., 2018], XQUAD [Artetxe et al., 2019], MLQA [Lewis et al., 2019] and Belebele [Bandarkar et al., 2023]. This parallel nature facilitates monolingual, multilingual, and cross-lingual experiments, enabling valuable comparisons across languages. However, this approach introduces biases related to translations and cultural representation, affecting experimental conclusions by reflecting the origin culture.
Recently, there has been a focus on creating native evaluation benchmarks to assess local cultural knowledge, rather than relying on translations from English. These native datasets, which resemble popular English benchmarks, include unique cultural elements that are generally more challenging for current models. They usually are of higher quality than machine or human-translated datasets. For example, native MMLU [Hendrycks et al., 2020] datasets have been created for Chinese [Li et al., 2023], Korean [Son et al., 2024], Indonesian [Koto et al., 2023] and Arabic [Koto et al., 2024]. Other examples of language-specific evaluation benchmarks include C-Eval for Chinese [Huang et al., 2024], HAE-RAE Bench for Korean [Son et al., 2023], COPAL-ID for Indonesian [Wibowo et al., 2023] and RoCulturaBench for Romanian [Masala et al., 2024]. Finally, Etxaniz et al. [2024] introduces 4 native Basque multiple-choice datasets with local questions.
@@ 6 Conclusion and Future Work
While our work is a first step in this direction, it also comes with some limitations, which we would like to address in future work. More concretely, the local subset of our benchmark is limited to questions about the Basque culture. While we expect that the general trends we observe would also apply to other minority cultures, we believe that it would be valuable to build similar datasets covering other cultures. This is generally more challenging than developing benchmarks about global subjects, as it usually requires being part of or engaging with the relevant communities. We hope that our work prompts to reconsider how LLMs are evaluated more broadly, and motivates the creation of similar datasets for other minority cultures. Besides, the English version of our dataset was created through professional translation, which could lead to translationese and other translation artifacts [Artetxe et al., 2020]. This typically occurs in the opposite direction, as most datasets are translated from English into other languages. In the future, we would like to analyze if this has any impact in our results, and explore authoring questions about minority cultures in English rather than translating from their respective languages.
@@ Acknowledgements
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Included in Section 6
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] We see no negative societal impacts
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] No training is done, only evaluations
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A] The evaluation is deterministic
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Included in Appendix D
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [No] We include references
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] There is no personally identifiable information or offensive content
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix E Results by Category
We show that previous local and global results are consistent across categories in Tables 10 and 11. The large difference between local and global is maintained across all models and categories. When comparing Llama 2 and Latxa, we see that previous results are consistent across all categories. That is, Latxa is better at local questions and Llama 2 is better at global questions. However, we see that the differences vary significantly depending on the models and categories.
@@ Appendix F Results by Difficulty
Results are also consistent across difficulty levels, as shown in Tables 10 and 13. All models obtain worse results in local questions at all difficulty levels. There is a clear difference between difficulties, with the biggest drop in performance happening from the easy to medium difficulty. Commercial models have a bigger drop in scores on local questions than on global questions for medium and hard questions. For open models, results are more varied, but the relative drop is also generally bigger on local questions. On the most difficult local questions, some models get close to random chance.
@@
Appendix G Extended Error Analysis
We include the remaining 25 examples of error analysis in Table 14.
Table 14: Last 25 error analysis examples annotated by difficulty of web search.
Category Diff. Question Candidate 0 Candidate 1 Candidate 2 Ans. Web Sports and Leisure 3 Against whom did the "aizkolari" or woodchopper Joxe Mari Olasagasti first compete? Against Floren Nazabal Against Donato Larretxea Against Jose Etxebeste 2 difficult Music and Dance 2 Whose song is "Freaky Fiesta"? The Skalariak group’s Betagarri’s Kortatu’s 1 easy Cinema and Shows 2 Who is the protagonist in Borja Cobeaga’s short film "The First Time"? Txema Blasco Marivi Bilbao Barbara Goenaga 1 difficult Cinema and Shows 1 In the show "Vaya Semanita" where is the character El Jonan from? From Otxarkoaga From Zorroza From Barakaldo 2 easy Music and Dance 3 Where exactly is the Biscayan group "Izenik ez" from? From Romo Fron Abadiño From Ugao-Miraballes 0 difficult Cinema and Shows 1 Which festivity in the Basque Country appears in the film "Day & Night"? The Aste Nagusia or Great Week of Bilbao The Tamborrada of Donostia / San Sebastian The San Fermín festival of Pamplona / Iruña 2 no Music and Dance 3 Who currently directs the Alurr group in Ibarra? Unax Sarriegi Aiert Beobide Edu Muruamendiaraz 1 no Sports and Leisure 2 Where’s the Onena leisure association from? From Irun From Ordizia From Donostia / San Sebastian 2 difficult Society and Traditions 3 Where is the oldest church in Biscay? In Lekeitio In Durango In Bilbao 1 no Music and Dance 2 Where’s the Pi LT band from? From the Mungia area From Amorebieta-Etxano From Bermeo 0 easy Science and Technology 2 When was Pamplona / Iruña General Hospital founded? In 1555 In 1545 In 1554 1 no Sports and Leisure 2 When is Navarre Day celebrated? The first Sunday in May The last Sunday in April The last Saturday in April 1 difficult Culture and Art 1 How are Sebastian and Joxe Lizaso related? They are brothers They are cousins They are father and son 2 easy Music and Dance 2 Where’s the San Fermín dance company from? From Zizur From Tafalla From Tudela 0 no Cinema and Shows 3 Where was the Basque screenwriter and director Pello Varela born? In Amorebieta-Etxano In Vitoria-Gasteiz In Portugalete 1 easy Basque and Literature 2 How often is the Jakin magazine published? Once a month From time to time Once a year 0 difficult Basque and Literature 3 When did the translation and interpretation degree course start at the University of the Basque Country (UPV/EHU)? In the 1995-1996 academic year In the 1990-1991 academic year In the 2000-2001 academic year 2 no Science and Technology 3 What’s the only amphibian on Mount Urgull? The common frog The salamander The common toad 1 difficult Geography and History 1 Where’s the Gorbeialdea? In Biscay In Álava and Biscay In Álava 2 easy Music and Dance 1 Where’s the Getaria dance company from? From Labourd From Álava From Gipuzkoa 0 easy Sports and Leisure 3 In 2010, which women’s team won the Gipuzkoa trainera rowing championship? San Juan Getaria-Tolosa Zumaia 2 no Society and Traditions 2 Which of the following is not the function of the chartered councils? Approval of the chartered regulations Promotion of cultural activities Tax collection 0 no Basque and Literature 3 What was the first novel in Basque to be published in the Southern Basque Country after the War? "Hiltzaileak" "Alos-Torrea" "Leturiaren egunkari ezkutua" 1 difficult Sports and Leisure 3 Where is the Xaguxar leisure group from? From Astigarraga From Hernani From Ordizia 0 easy Music and Dance 3 Who was the first chairman of the Euskal Dantzarien Biltzarra (Association of Basque Dancers)? Jesus Maria Arozamena Jose Antonio Legarra Juan Jose Garaizabal 0 difficult